### v1

@@ Abstract
Significant progress has been made in the field of handwritten mathematical expression recognition, while existing encoder-decoder methods are usually difficult to model global information in . Therefore, this paper introduces a novel approach, Implicit Character-Aided Learning (ICAL), to mine the global expression information and enhance handwritten mathematical expression recognition. Specifically, we propose the Implicit Character Construction Module (ICCM) to predict implicit character sequences and use a Fusion Module to merge the outputs of the ICCM and the decoder, thereby producing corrected predictions. By modeling and utilizing implicit character information, ICAL achieves a more accurate and context-aware interpretation of handwritten mathematical expressions. Experimental results demonstrate that ICAL notably surpasses the state-of-the-art(SOTA) models, improving the expression recognition rate (ExpRate) by 2.21%/1.75%/1.28% on the CROHME 2014/2016/2019 datasets respectively, and achieves a remarkable 69.25% on the challenging HME100k test set. We make our code available on the GitHub. [^1] https://github.com/qingzhenduyu/ICAL
@@ 1 Introduction
  • Experimental results indicate that the ICAL method surpasses previous state-of-the-art methods and achieves expression recognition rate (ExpRate) of 60.59%, 58.73%, and 60.40% on the CROHME 2014 [14]/2016 [15]/2019 [13] test sets, respectively, and an ExpRate of 69.25% on the HME100K test set [28].
@@ 4.4 Comparison with State-of-the-art Methods
Table  1 presents the results on the CROHME dataset. To ensure fairness in performance comparison and considering that different methods have used various data augmentation techniques, many of which have not been disclosed, we have limited our comparison to results without the application of data augmentation.
@@
As shown in table 1, our method achieved the best performance across all metrics. Our method outperforms CoMER by 2.21%/1.75%/1.28% on the CROHME 2014, 2016, and 2019 datasets, respectively. Across all metrics, the ICAL method achieves an average improvement of 1.6% over the CoMER method. The experimental results on CROHME dataset prove the effectiveness of our method.
We also conducted experiments on the challenging and real-world dataset HME100K, as shown in table 2. We have replicated the performance of CoMER on the HME100K dataset, and our method surpasses the state-of-the-art(SOTA) by 1.13%, reaching an impressive 69.25%. The outstanding experimental performance on the HME100K dataset, which is more complex, larger, and more realistic compared to CROHME, further proves the superior generalization ability and effectiveness of the ICAL method.
@@ 5 Conclusion
In this paper, we propose a novel recognizer framework, ICAL, capable of leveraging global information in  to correct the predictions of the decoder. Our main contributions are threefold: (1) We have designed an Implicit Character Construction Module(ICCM) to predict implicit characters in . (2) Additionally, we employ a Fusion Module to aggregate global information from implicit characters, thereby refining the predictions of the Transformer Decoder. We integrate these two modules into the CoMER model to develop our method, ICAL. (3) Experimental results demonstrate that the ICAL method surpasses previous state-of-the-art approaches, achieving expression recognition rate (ExpRate) of 60.59%, 58.73%, and 60.40% on the CROHME 2014, 2016, and 2019 datasets, respectively, and an ExpRate of 69.25% on the HME100K dataset.

### v2

@@ Abstract
Significant progress has been made in the field of handwritten mathematical expression recognition, while existing encoder-decoder methods are usually difficult to model global information in . Therefore, this paper introduces a novel approach, Implicit Character-Aided Learning (ICAL), to mine the global expression information and enhance handwritten mathematical expression recognition. Specifically, we propose the Implicit Character Construction Module (ICCM) to predict implicit character sequences and use a Fusion Module to merge the outputs of the ICCM and the decoder, thereby producing corrected predictions. By modeling and utilizing implicit character information, ICAL achieves a more accurate and context-aware interpretation of handwritten mathematical expressions. Experimental results demonstrate that ICAL notably surpasses the state-of-the-art(SOTA) models, improving the expression recognition rate (ExpRate) by 2.25%/1.81%/1.39% on the CROHME 2014/2016/2019 datasets respectively, and achieves a remarkable 69.06% on the challenging HME100k test set. We make our code available on the GitHub. [^1] https://github.com/qingzhenduyu/ICAL
@@ 1 Introduction
  • Experimental results indicate that the ICAL method surpasses previous state-of-the-art methods and achieves expression recognition rate (ExpRate) of 60.59%, 58.73%, and 60.40% on the CROHME 2014 [14]/2016 [15]/2019 [13] test sets, respectively, and an ExpRate of 69.06% on the HME100K test set [28].
@@ 4.4 Comparison with State-of-the-art Methods
Table  1 presents the results on the CROHME dataset. To ensure fairness in performance comparison and considering that different methods have used various data augmentation techniques, many of which have not been disclosed, we have limited our comparison to results without the application of data augmentation. Given the relative small size of the CROHME dataset, we conducted experiments with both the baseline CoMER and the proposed ICAL model using five different random seeds (7, 77, 777, $7$, 77777) under the same experimental conditions. The reported results are the averages and standard deviations of these five experiments.
@@
As shown in table 1, our method achieved the best performance across all metrics. Our method outperforms CoMER by 2.25%/1.81%/1.39% on the CROHME 2014, 2016, and 2019 datasets, respectively. Across all metrics, the ICAL method achieves an average improvement of 1.6% over the CoMER method. The experimental results on CROHME dataset prove the effectiveness of our method.
We also conducted experiments on the challenging and real-world dataset HME100K, as shown in table 2. We have replicated the performance of CoMER on the HME100K dataset, and our method surpasses the state-of-the-art(SOTA) by 0.94%, reaching an impressive 69.06%. The outstanding experimental performance on the HME100K dataset, which is more complex, larger, and more realistic compared to CROHME, further proves the superior generalization ability and effectiveness of the ICAL method.
@@ 5 Conclusion
In this paper, we propose a novel recognizer framework, ICAL, capable of leveraging global information in  to correct the predictions of the decoder. Our main contributions are threefold: (1) We have designed an Implicit Character Construction Module(ICCM) to predict implicit characters in . (2) Additionally, we employ a Fusion Module to aggregate global information from implicit characters, thereby refining the predictions of the Transformer Decoder. We integrate these two modules into the CoMER model to develop our method, ICAL. (3) Experimental results demonstrate that the ICAL method surpasses previous state-of-the-art approaches, achieving expression recognition rate (ExpRate) of 60.59%, 58.73%, and 60.40% on the CROHME 2014, 2016, and 2019 datasets, respectively, and an ExpRate of 69.06% on the HME100K dataset.